If the mirror separation d = 50 cm changes by 2 nm, what would be the change in relative frequency? If the laser frequency is νL = 4 x 1014s-1, what would be the value of ∆νL=?
1
Expert's answer
2021-01-14T10:37:24-0500

Answer

Relative frequency is given

"\lambda=0.5m"

"\Delta \lambda=2nm"

Frequency

"\nu=\frac{c}{\lambda}=6\times10^8Hz"

Now relative frequency is given

"\frac{\Delta \nu}{\nu}=\frac{c\Delta \lambda}{\lambda^2}=\frac{3\times10^8\times2\times10^{-9}}{0.5\times0.5}"

"=12"

"\Delta \nu_L=12\times4 \times 10^{14}\\=48 \times 10^{14}Hz"